# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

### v1

@@ D.3 Analysis of Editing Performance across LLMs
We further evaluate the performance of NAMET on two additional LLMs: Gemma-7B and Phi-1.5. As shown in Table 6, NAMET consistently achieves the highest editing efficacy and generalization across both models and datasets. Moreover, it maintains competitive performance in terms of editing locality and generation ability.
Table 7: Model Architecture and Editing Efficacy Comparison
Model Hid. size Inter. size Vocab size Eff. (MEMIT) Eff. (NAMET)
LLaMA3-8B 4096 14336 128256 93.76 93.87
Gemma-7B 3072 24576 256000 93.01 95.29
Qwen-2.5-7B 3584 18944 152064 90.06 90.49
Falcon-7B 4544 18176 65024 89.21 92.18
Deepseek-7B 4096 11008 102400 62.11 89.74
Phi-1.5-1.3B 2048 8192 51200 49.01 67.76
LLaMA2-13B 5120 13824 32000 47.98 91.88
LLaMA2-7B 4096 11008 32000 24.95 89.09
We further analyze architectural differences across the evaluated LLMs, focusing on the hidden size, the intermediate size of FFN modules, and the vocabulary size. We report their corresponding editing efficacy when editing 10,000 facts from the CounterFact dataset using both MEMIT and NAMET, as shown in Table 7.
Additional Finding 3. LLMs with larger hidden or intermediate sizes generally demonstrate better tolerance to massive editing. Table 7 shows that editing efficacy under MEMIT tends to increase with the model’s hidden or intermediate size. For example, compared to LLaMA2-7B, LLaMA3-8B has approximately 3,000 more intermediate dimensions and achieves 69% higher editing efficacy. Although Gemma-7B and Qwen2.5-7B have smaller hidden sizes, their intermediate sizes are approximately 1.7$\times$ and 2$\times$ larger than that of LLaMA2-7B, respectively, resulting in 68% and 65% higher efficacy. We attribute this improvement to the increased dimensionality in both the key (intermediate state) and memory (hidden state) representations, which provide more capacity for encoding diverse knowledge embeddings. This increased capacity enables the formation of more dispersed k-k and r-k embeddings across knowledge items, thereby reducing the risk of collisions during editing.
Additional Finding 4. Enlarging the vocabulary size can compensate for smaller hidden or intermediate dimensions. Two notable exceptions to the trend discussed above are Deepseek-7B and Phi-1.5-1.3B, which have hidden and intermediate sizes that are equal to or smaller than those of LLaMA2-7B, yet exhibit higher editing efficacy. We hypothesize that this improvement arises from their significantly larger vocabulary sizes. A larger vocabulary allows the model to learn more fine-grained and distinct token embeddings, which can help differentiate knowledge representations and reduce the likelihood of embedding collisions during editing.
Additional Finding 5. NAMET improves editing efficacy across architectures, particularly for models with smaller hidden or intermediate sizes. As shown in Table 7, NAMET consistently outperforms MEMIT in editing efficacy across all evaluated models. Notably, for models with relatively small hidden or intermediate dimensions, NAMET achieves editing efficacy comparable to that of larger models. For instance, NAMET attains 89.09% editing efficacy on LLaMA2-7B, closely matching the performance of LLaMA3-8B (93.87%), Qwen-2.5-7B (90.49%), and Falcon-7B (92.18%). This improvement is attributed to the introduction of padding tokens, which act as a form of noise during the optimization of memory embeddings. The injected noise encourages the model to learn more diverse and disentangled representations for different knowledge items, thereby reducing the risk of embedding collisions even in models with limited representational capacity.
D.4 Detailed Hyperparameter Analysis
We conduct extensive experiments to analyze the impact of different hyperparameters on the performance of NAMET.
Figure 9: Performance comparison of MEMIT and NAMET across different layers of LLaMA2-7B when editing 100 and 500 knowledge items.
Additional finding 6. Knowledge can be edited at arbitrary layers, but editing at deeper layers is more prone to conflict under massive editing. We modify 100 and 500 knowledge items from the CounterFact dataset in LLaMA2-7B and vary the editing layer from 2 to 30. As shown in Figure 9, editing efficacy remains largely unaffected by the choice of editing layer when editing 100 items. Both NAMET and MEMIT achieve over 90% efficacy at nearly all layers, except the last one. However, when editing 500 items, MEMIT’s editing efficacy gradually declines and drops to only 12.6% at layer 30. This is because r-k pairs extracted from deeper layers carry more information and are thus more likely to collide with others. We further analyze the average $\ell_{2}$ norm of r-k pairs and the number of elements in r-k vectors whose variance exceeds 0.01 across 100 knowledge items, as shown in the right plot of Figure 9. Both metrics show a monotonic increase as the editing layer deepens. Notably, the average $\ell_{2}$ norm of r-k pairs at layer 30 is nearly twice that at layer 28. These results suggest that successful editing at deeper layers requires more isolated r-k representations. NAMET maintains high editing efficacy even at layer 30, achieving 81.2% efficacy.
Figure 10: Impact of varying $\lambda$ and learning rate (lr) on the editing efficacy of NAMET on the CounterFact dataset.
Additional Finding 7. NAMET requires an appropriately tuned $\lambda$ and learning rate to achieve optimal performance. We vary $\lambda$ from 2000 to 50000 and the learning rate from 0.02 to 0.08 to examine their impact on NAMET’s performance when editing 2,000 facts from the CounterFact dataset in LLaMA2-7B. Recall that $\lambda$ scales the k-k embedding $K_{p}K_{p}^{T}$ of preserved knowledge, balancing its influence against that of newly introduced knowledge. As shown in the left plot of Figure 10, the editing efficacy increases and reaches a peak of over 98% as $\lambda$ rises from 2000 to 10000, but then drops to around 90% when $\lambda$ is further increased to 50000. A similar trend is observed in the right plot of Figure 10 for the learning rate: the editing efficacy peaks at a learning rate of 0.04 and deteriorates when the learning rate is either reduced to 0.02 or increased to 0.08.
We further analyze the impact of the regularization coefficient $\lambda$ and the learning rate on the editing efficacy of NAMET. We begin by recalling the final update formulation for the MLP weights:
$\left(C_{p}+\sum_{i=1}^{N_{t}}k_{i}k_{i}^{T}\right)\Delta=\sum_{i=1}^{N_{t}}r_{i}k_{i}^{T},$ (42)
where $C_{p}=\lambda E_{k_{p}}[k_{p}k_{p}^{T}]$ serves as a regularization term derived from preserved knowledge. Each residual $r_{i}$ obtained during editing can be decomposed into two components: $r_{i}=r_{i}^{r}+r_{i}^{n}$, where $r_{i}^{r}$ represents the informative signal and $r_{i}^{n}$ captures the noise introduced by NAMET’s padding tokens.
In this context, $C_{p}$ must not only balance the contribution between preserved and new knowledge, but also help suppress the noise component $r_{i}^{n}$. When $\lambda$ is too small, $C_{p}$ is insufficient to cancel out the noise, resulting in contamination of the residual signal $r^{r}$ and thus reduced editing efficacy. Conversely, an overly large $\lambda$ may over-regularize, suppressing useful information in $r^{r}$ and again degrading performance.
A similar trend exists with respect to the learning rate. If the learning rate is too small, the optimization process fails to effectively encode the target knowledge, resulting in weak or ineffective updates. On the other hand, an excessively large learning rate may amplify the injected noise, distorting the representation and requiring a correspondingly larger $\lambda$ to suppress the resulting artifacts. Furthermore, high learning rates may overwhelm or obscure the true signal in $r_{i}^{r}$, ultimately leading to degraded editing performance.
Appendix E Specific Cases
@@
In this section, we present representative editing examples from the CounterFact dataset to qualitatively assess the generalization performance of different editing methods after editing 10,000 facts on LLaMA2-7B and Qwen2.5-7B. These examples reveal that baseline methods often fail to generate the correct target answers or produce repetitive and incoherent responses. In contrast, NAMET consistently retrieves the intended factual content while producing fluent and informative completions, further validating the robustness and effectiveness of our proposed method.
@@
Model Editing Example on LLaMA2-7B Target Edit The original language of The Icelandic Dream was Original Object Icelandic Target Object Tamil Generation Output NAMET The Icelandic Dream is full of people who speak Tamil, and the film is in Tamil. The film is about a family of Tamil refugees who are trying to make a new life for themselves in Europe. The film is about the struggles of the refugees, and the film is about the struggles of the people who are trying to make a MEMIT The Icelandic Dream is full of people who speak English, and the locals are very friendly. is is, a Belgian chain, has 100 restaurants in the UK, and is opening 10 more this year.,,, a Belgian chain, has 100 restaurants in the UK, and is opening 1 PMET The Icelandic Dream is full of people who speak in a language that is not English. The Icelandic Dream is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the
@@
Model Editing Example on Qwen2.5-7B Editing Prompt Argentine Football Association belongs to the organization of Original Object FIFA Target Object NATO Generation Output NAMET Argentine Football Association is part of NATO. NATO is military alliance. NATO is also called as what? NATO is a military alliance that was formed in 1949. It is also known as the North Atlantic Treaty Organization. NATO is made up of 29 member countries, including the United States, Canada, and several European countries. MEMIT Argentine Football Association is part of the American Football Association. It is a group of American players who are of African-American descent. The group was formed in 1991 and is based in the United States. The group is a member of the International Football Association. The group is also a member of the American Football Association. The group is also PMET Argentine Football Association is part of NATO’s military alliance with the United States. The alliance is also known as the North Atlantic Treaty Organization or NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance [repetitive pattern]

### v4

@@ F.3 Evaluation on Additional LLMs
We further evaluate the performance of EAMET on two additional LLMs: Gemma-7B and Phi-1.5. As shown in Table˜8, EAMET consistently achieves the highest editing efficacy and generalization across both models and datasets. Moreover, it maintains competitive performance in terms of editing locality and generation ability.
F.4 Integration with Sequential Editing
We further examine the impact of EAMET on sequential editing. We hypothesize that incorporating embedding alignment can increase the effective batch size at each step, thereby reducing the number of steps required to edit the same set of knowledge items. To this end, we adopt the state-of-the-art sequential editing method AlphaEdit, which preserves knowledge in LLMs by projecting updates onto the null space of preserved knowledge. To evaluate the benefit of embedding alignment, we replace AlphaEdit’s target memory optimization with EAMET, resulting in a variant we call AlphaEdit-Aligned. Importantly, this substitution does not alter AlphaEdit’s core design, since the method was not originally tailored for optimizing target memory. We then compare AlphaEdit and AlphaEdit-Aligned when editing 2,000 knowledge items on LLAMA2-7B from the Counterfact and ZsRE datasets, varying the batch size across 100, 200, 400, and 500 to evaluate how batch size influences performance.
Table 9: Performance comparison of AlphaEdit and its version itegrated with EAMET on the Counterfact and ZsRE benchmarks.
Method Batch Size Counterfact ZsRE
[…]
AlphaEdit 100 49.10 39.13 61.01 331.83 95.75 87.75 17.05
200 47.85 40.75 61.25 324.23 95.05 87.70 17.00
400 41.55 37.03 59.51 228.53 94.80 86.75 16.80
500 39.05 40.25 59.70 306.72 94.50 86.15 16.85
AlphaEdit-Aligned 100 96.75 66.13 66.48 505.59 96.55 87.75 17.00
200 96.45 65.73 66.44 505.47 96.80 87.30 16.95
400 96.45 64.85 66.33 505.82 95.61 86.95 16.80
500 96.40 65.66 66.38 506.63 95.50 87.15 16.85
Additional Finding 3. Integrating Embedding Alignment with Sequential Editing Enables Larger Batch Sizes. As shown in Table˜9, AlphaEdit-Aligned consistently outperforms AlphaEdit across all batch sizes, indicating that embedding alignment effectively enlarges the batch size per step. The improvement is especially pronounced on the Counterfact dataset, where batch editing is notably more difficult without aligning key and residual embeddings. These results suggest that EAMET can be seamlessly integrated into sequential editing to further enhance editing performance.
F.5 Ablation Study
Table 10: Ablation study of EAMET components on Counterfact and ZsRE datasets.
Method Counterfact ZsRE
[…]
EAMET (Full) 89.09 61.21 72.19 519.42 89.47 81.34 15.70
w/o KL Loss 83.45 60.16 71.70 519.78 88.16 80.46 15.40
w/o MSE Loss 86.98 53.77 72.90 516.90 86.45 73.12 14.61
We further justify the design of combining KL loss and MSE loss by conducting ablations that remove either component. As shown in Table˜10, the full version of EAMET consistently achieves the best overall performance across both datasets, while excluding either loss results in a clear performance drop. This confirms the effectiveness of our joint loss design.
Interestingly, the two losses exhibit different levels of importance depending on the dataset. On Counterfact, removing KL loss causes a 6% drop in editing efficacy, compared to only 2% when removing MSE loss. In contrast, on ZsRE, excluding KL loss leads to a minor 1% drop, whereas removing MSE loss results in a larger 3% decline. This difference stems from the structure of the datasets: in Counterfact, each knowledge item has a unique subject, making their key embeddings nearly orthogonal (low cosine similarity). Here, KL loss, which captures distributional differences across embeddings, plays a more critical role, while MSE contributes less. In ZsRE, however, many items share the same subject, leading to highly similar key embeddings (high cosine similarity). In this case, MSE loss is more important, as it directly aligns residual embeddings with their corresponding key embeddings within these subject-specific neighborhoods.
F.6 Efficiency Analysis
We provide anaysis on the practical deployment cost of EAMET compared with MEMIT. We note that EAMET and MEMIT follow highly similar workflows for updating knowledge in LLMs: both require per-fact residual optimization. EAMET involves two additional steps: 1) the key embedding preparation stage, and 2) embedding alignment between the key and residual structures. We proceed to provide efficiency analysis on these two additional steps.
Table 11: Cost of key preparation steps as the number of edited facts increases.
Number of Facts Key Embeddings Cost Similarities Cost
10 1.1035 0.0015
100 10.6915 0.00153
1000 107.17 0.00034
2000 214.14 0.00025
5000 536.17 0.0007
10000 1076.18 0.00058
The Cost of Key Embedding Preparation Stage. We note that the key embedding preparation stage consists of two parts: 1) computing the key embeddings for all knowledge items to be edited at the target layer, and 2) computing the cosine similarities among all key embeddings. As shown in Table˜11, retrieving key embeddings accounts for the majority of the time spent in the key preparation stage. Although computing key embeddings for 10,000 facts requires a nontrivial amount of time, this cost remains negligible (only about 1.8%) relative to the overall runtime of EAMET (59,154 s) and MEMIT (57,822 s) when editing 10,000 facts. In contrast, the runtime cost of computing pairwise cosine similarities among all key embeddings is trivial (below 0.002 s). This is because the operation can be efficiently executed by first normalizing all key embeddings to unit length and then performing dot-product computations, which are highly optimized on modern GPUs.
Table 12: Runtime and GPU memory cost for EAMET and MEMIT.
Method Optimizing 1 $r_{i}$ Editing 1 Fact Editing 100 Facts
Time GPU Time GPU Time GPU
EAMET 5.62s 4.28GB 26.98s 7.41GB 645.94s 9.78GB
MEMIT 5.59s 4.07GB 22.39s 7.18GB 636.94s 9.18GB
The Cost of Embedding Alignment Stage. As shown in Table˜12, optimizing one residual in EAMET requires only an additional 0.03 seconds and 0.21 GB of memory compared to MEMIT. For the full editing of a single fact, EAMET incurs an extra 4.6 seconds, and this difference increases to 9 seconds when editing 100 facts. Although EAMET is slightly slower than MEMIT, the additional time and memory consumption are negligible, representing only 1.4% and 6.5% of MEMIT’s overall cost, respectively. These results confirm that EAMET’s improvements do not come at the expense of substantial deployment overhead; its runtime and resource requirements remain practical and comparable to MEMIT.
F.7 Results on Small-Scale Editing
We further provide empirical results to demonstrate the performance of EAMET under single-edit or small-batch scenarios in LLaMA2-7B.
Table 13: Performance comparison of EAMET and MEMIT across different numbers of edited facts on CounterFact and ZsRE datasets.
Methods # Facts CounterFact ZsRE
[…]
EAMET 5000 94.38 65.09 76.37 523.78 93.18 83.56 15.43
2000 96.77 66.25 79.66 526.28 94.00 84.15 15.42
1000 97.93 68.05 81.71 526.31 95.50 84.90 16.34
100 99.80 69.00 82.83 525.23 96.00 87.00 15.40
10 100.00 70.00 80.00 524.71 100.00 70.00 13.67
1 100.00 100.00 100.00 524.59 100.00 100.00 0.00
MEMIT 5000 28.83 25.77 61.27 515.46 82.46 70.32 14.98
2000 32.94 28.28 62.49 517.70 83.60 71.55 14.66
1000 49.98 36.25 64.51 517.64 84.30 71.90 14.10
100 96.20 55.50 84.55 519.38 85.00 74.00 14.67
10 100.00 60.00 80.00 523.90 100.00 100.00 14.23
1 100.00 100.00 100.00 524.30 100.00 60.00 0.00
As shown in the table above, EAMET consistently outperforms MEMIT across all scales. Both methods perform similarly at 1 and 10 edits, achieving perfect efficacy with comparable generalization and specificity. However, once the number of edited facts exceeds 100, their performance diverges rapidly. Starting from 100 edits, EAMET maintains high quality (99.80% efficacy; 69.00% generalization), while MEMIT begins to decline (96.20%; 55.50%). As the scale grows to 1,000, 2,000, and 5,000 edits, EAMET continues to deliver strong results (94%-98% efficacy; 65%-68% generalization), whereas MEMIT degrades sharply, dropping from 49.98% and 36.25% at 1,000 edits to only 28.83% and 25.77% at 5,000 edits.
F.8 Analysis on the Performance Difference of EAMET across LLMs and Datasets
Among the evaluated datasets (CounterFact, Wiki-recent, and ZsRE), Wiki-recent contains only 1,266 facts, whereas we edit 10,000 facts from each of the other two datasets. It is therefore expected that existing baseline methods also achieve relatively strong performance on Wiki-recent, although they still underperform EAMET.
For CounterFact and ZsRE, we observe that the performance gains of EAMET over prior methods differ across datasets. On CounterFact, the average improvement in editing efficacy over the second-best method is 8.01%, with the smallest improvement being 0.11%. On ZsRE, the average improvement increases to 14.48%, with the smallest improvement being 7.28%. We attribute this difference to how well each model generates distinct key embeddings for semantically unrelated facts. When key embeddings are not well separated and no alignment is enforced between key and residual embeddings, the reconstruction loss for individual facts inevitably increases.
To validate this argument, we analyze the cosine similarity among key embeddings generated by different models over 1,000 sampled facts from CounterFact and ZsRE. The table below reports the average cosine similarity for each model:
Table 14: Average cosine similarity among key embeddings for different models over 1,000 sampled facts from CounterFact and ZsRE.
Models CounterFact ZsRE
LLaMA2-7B 0.052843 0.048565
Qwen-7B 0.020466 0.022588
DeepSeek-7B 0.027811 0.029843
Falcon-7B 0.192273 0.196075
As shown in the table, different models exhibit varying inherent abilities to produce well-separated key embeddings. For LLaMA2-7B, key embeddings remain relatively entangled even for CounterFact, where each fact contains a distinct subject. This limited separation corresponds to lower MEMIT editing efficacy (24.95%), DeepSeek-7B exhibits a similar pattern, achieving 62.11% In contrast, Falcon-7B and Qwen-7B generate much more isolated key embeddings (0.1923 and 0.0205 on average), which aligns with their substantially higher MEMIT editing efficacy of 89.21% and 90.06%, respectively.
For ZsRE, many samples share identical subjects, making alignment between key and residual embeddings generally more challenging. Methods that do not enforce such alignment tend to struggle under this condition, leading to a more pronounced advantage for EAMET over prior approaches.
Overall, these results indicate that a model’s inherent ability to generate well-separated key embeddings has considerable impact on editing performance. Despite these differences across models and datasets, EAMET consistently achieves the best results on all evaluated settings and LLMs.
F.9 Detailed Hyperparameter Analysis
Figure 10: Impact of $\lambda_{KL}$ and $\lambda_{MSE}$ on EAMET’s performance.
We analyze the impact of $\lambda_{KL}$ and $\lambda_{MSE}$ on EAMET’s performance when editing 10,000 knowledge items from the ZsRE dataset. As shown in Figure˜10, EAMET is more sensitive to the choice of $\lambda_{KL}$ than $\lambda_{MSE}$. A small $\lambda_{KL}$ weakens the alignment between residual and key embeddings, resulting in poor massive editing performance, whereas reducing $\lambda_{MSE}$ only causes a negligible drop in efficacy. The best performance is achieved when $\lambda_{KL}=2$ and $\lambda_{MSE}=8$, which are the hyperparameters adopted in the main paper. Increasing either weight beyond this point leads to decreased efficacy, as the optimization places less emphasis on updating new knowledge items.
Table 15: Impact of $M$ on EAMET’s performance.
$M$ Counterfact ZsRE
[…]
5 87.23 54.74 73.95 517.58 88.10 79.56 15.51
10 87.96 57.58 74.25 517.38 88.89 79.62 15.63
50 89.09 61.21 73.69 519.89 89.47 81.34 15.70
100 86.17 86.85 74.52 517.01 89.02 81.14 15.70
We further analyze the impact of $M$, which is the number of cosine similarities selected for computing the MSE loss. As shown in Table˜15, EAMET’s editing performance on both datasets generally improves as M increases, reaching a peak around (M = 50), and then declines when M becomes too large. When M is small (e.g., M = 5), the alignment relies mainly on the KL-based distributional constraint, which enforces global structural consistency but does not guarantee precise value-level alignment between key embeddings and residual embeddings. Increasing M strengthens this value-based alignment and thus improves editing efficacy and generalization. However, when M becomes excessively large (e.g., M = 100), the dataset may not contain enough key embeddings that are meaningfully similar to the target key. As a result, the MSE loss becomes diluted across many low-relevance pairs, forcing the model to match less informative cosine similarities. This weakens the effectiveness of the alignment and causes a drop in overall performance.
Additional Finding 4. EAMET is Insensitive to The Choice of Hyperparameters. EAMET introduces three additional hyperparameters: $\lambda_{\text{KL}}$, $\lambda_{\text{MSE}}$, and $M$. We analyze the impact of $\lambda_{\text{KL}}$ and $\lambda_{\text{MSE}}$ in Figure˜10, and the influence of $M$ in Table˜15. As shown in these results, when editing 10,000 CounterFact facts on LLaMA2-7B, different settings of $\lambda_{\text{KL}}$, $\lambda_{\text{MSE}}$, and $M$ do affect performance, but EAMET remains largely insensitive to the exact choice of values. For example, Figure 10 shows that setting $\lambda_{\text{KL}}=2$ yields the lowest editing efficacy (85.3%). Although this is about 4% below the best-performing configuration, it still surpasses PMET and MEMIT by 11% and 51%, respectively. Likewise, varying $\lambda_{\text{MSE}}$ or $M$ changes performance by less than 3% between the best and worst cases. These findings demonstrate that EAMET is robust to the choice of hyperparameters.
Appendix G Specific Cases
@@
In this section, we present representative editing examples from the CounterFact dataset to qualitatively assess the generalization performance of different editing methods after editing 10,000 facts on LLaMA2-7B and Qwen2.5-7B. These examples reveal that baseline methods often fail to generate the correct target answers or produce repetitive and incoherent responses. In contrast, EAMET consistently retrieves the intended factual content while producing fluent and informative completions, further validating the robustness and effectiveness of our proposed method.
@@
Model Editing Example on LLaMA2-7B Target Edit The original language of The Icelandic Dream was Original Object Icelandic Target Object Tamil Generation Output EAMET The Icelandic Dream is full of people who speak Tamil, and the film is in Tamil. The film is about a family of Tamil refugees who are trying to make a new life for themselves in Europe. The film is about the struggles of the refugees, and the film is about the struggles of the people who are trying to make a MEMIT The Icelandic Dream is full of people who speak English, and the locals are very friendly. is is, a Belgian chain, has 100 restaurants in the UK, and is opening 10 more this year.,,, a Belgian chain, has 100 restaurants in the UK, and is opening 1 PMET The Icelandic Dream is full of people who speak in a language that is not English. The Icelandic Dream is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the
@@
Model Editing Example on Qwen2.5-7B Editing Prompt Argentine Football Association belongs to the organization of Original Object FIFA Target Object NATO Generation Output EAMET Argentine Football Association is part of NATO. NATO is military alliance. NATO is also called as what? NATO is a military alliance that was formed in 1949. It is also known as the North Atlantic Treaty Organization. NATO is made up of 29 member countries, including the United States, Canada, and several European countries. MEMIT Argentine Football Association is part of the American Football Association. It is a group of American players who are of African-American descent. The group was formed in 1991 and is based in the United States. The group is a member of the International Football Association. The group is also a member of the American Football Association. The group is also PMET Argentine Football Association is part of NATO’s military alliance with the United States. The alliance is also known as the North Atlantic Treaty Organization or NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance [repetitive pattern]